D.Thebonds were mind E The market issued 10-yar, 8% bonds witha value of S200,000. The company nds. Using the straight-line method, the amount of interest expense for the S1RL000 for the jest somiannual interest period is: (2 points) A 58.00000 B$8 500.00 C. S16,000.00 D $7,500.00 E $18,000.00 Os Jasary 1, 2010 d an 58,000, %, S yeur bond payable and is puyable On January 1, 2014, Jacob issues S800.000 of 9%, 13-year bonds at a price of 96%. What is the jounal entry to record the issuance of the bonds on January 1, 2014? (2 points) 800,000 800,000 A. Bonds Payable 800,000 800,000 B L Cash 800,000 Bonds Payable .Discount on Bonds Payable 772,000 28,000 Cash on Bonds Payable 28,000 800,000 D. Bonds Payable 772,000 28,000 Discount on Bonds Payable Bonds Payable E. 800,000
Answer #1

37. Answer is option B $103500

(100000*1.035) = $103500

38. Answer is option C

Cash = (850000*6%/2)= 25500

Discount on Bonds payable = (850000-825000)/10=2500

39. Answer is option B $8500

Cash = (200000*8%/2)=8000

Discount on Bonds payable = (200000-190000)/20=500

Total interest expense = 500

40. Answer is option E

Cash (800000*96.5%) Dr. 772000

Discount on Bonds payable (800000-772000)Dr. 28000

Bonds payable cr. 800000